A chemist needs 12 liters of 50% solution for a lab experiment. The stockroomconly has 40% and 70% solution. How much of each should be mixed together to make 12 liters of a 50% solution?
Answer by addingup(3677)

x+y= 12 and if we want to know just x, we subtract y:
x+y-y= 12-y
x= 12-y We'll use this value for x next.
---------------------------------------------------------
.40x+.70y= .50(12) Substitute for x:
.40(12-y)+.70y= .50(12)
4.8-.40y+.70y= 6 Subtract 4.8 on both sides and add/subtract y on left:
.30y= 1.2 Divide both sides by .30:
y= 4 He needs 4 liters of the 70% solution and:
12-4= 8 liters of the 40% solution
J